What is the number of significant figures and the result of:

485/9.231

12.62+1.5+025

0.071+4.36

please help? show work so i can understand how to do it? thanks!
1
Expert's answer
2011-09-02T07:12:30-0400
485 / 9.231 = 52.5403532 ≈ 52.54
12.62+1.5+0.25 = 14.12 + 0.25 = 14.37
12.62
+ 1.5
+ 0.25
14.37

0.071+4.36 = 4.431
0.071

+4.36
4.431